How pressure depends on fermi energy write only expression. For photon
1
Expert's answer
2020-11-18T06:33:39-0500

Fermi gas is a system of fermions, so, most likely, you meant proton, not photon, since photons are bosons. The equation is


"P=-\frac{\partial E_T}{\partial V}=\frac{2}{5}\frac NVE_F=\frac{(3\pi^2)^{2/3}\hbar^2}{5m_p}\bigg(\frac NV\bigg)^{5/3}"